how much would i need to invest now if i need 25,000 in 6 years time and i can invest @ 6% simple interest?
Answer by Theo(13342)

Simple interest is not compounded.

The formula would be:

y = x + (.06x*6) = x + .36x = 1.36x

You have y = 1.36*x

Divide both sides of this equation by 1.36 to get x = y/1.36

If y = 25000, then x = 25000/1.36 = 18382.35294

You have to invest 18382.35294 today to have 25000 in 6 years at 6% per year simple interest (no compounding).